# How to Beat Nakamoto in the Race

### v1

@@ 4.6. State Comparison Rules
(25) $[m,d,n,(l_{d\land m},\dots,l_{d\land m^{\prime}},\dots,l_{d}),I]\preceq[m^{\prime},d,n^{\prime},(l_{d\land m^{\prime}},\dots,l_{d}),I].$
@@
Equation (24) is true by Lemma 5. Higher lower and higher branches, $M,N$ improve the probability of satisfying (23), which leads to (25) . ∎
@@
We can classify states based on the timer of the height just above the lower branch, $l_{m+1}$. Reaching $\mathscr{A}$ requires the lower branch to be no lower than the public height. The value of $l_{m+1}$ provides information about the race between the rate of the jumper’s height increase (or public height increase) and the lower branch’s height increase.
@@
Our categorization of states bears resemblance to the “cold”, “critical” and “hot” regions defined in (gazi2020tight, ) in the sense that both reflect the adversary’s current advantage. The state categorization of (gazi2020tight, ) is based on the entire blocktree and all honest views, specifically the length difference between the adversary’s best chain and the longest honest chain visible to all honest nodes. In contrast, our state categorization is based on a simplified, sufficient state representation, determined by the timer just above the second-highest branch, which allows a precise analysis of the MDP.
@@ 5. The Zero-Delay Special Case
This section is devoted to the special case of $\Delta=0$, where all honest nodes are fully synchronized at all times. An important consequence is that all H-blocks are on different heights. Also, all timers are either zero or infinity, so the public height $P=d$. The sequence of timers in the representation (21) always takes the form of a sequence of $\max\{1,d-m+1\}$ zeros. In this section, we omit this sequence of timers to use $[m,d,n,I]$ as a sufficient representation of the state, which is further reduced to $[m,d,n]$ when $I=\emptyset$.
@@ 5.1. Preliminaries
The state $[m,d,n,H]$ includes a new H-block arrival, which can only be added above the public height, $d$. Specifically, it must be added to either the higher branch on a height between $d+1$ and $n+1$ or the lower branch on a height between $d+1$ and $m+1$ (this is applicable only if $d\leq m$). If the block is added to the higher branch on some height $i\in\{d+2,\dots,n\}$, then the system transitions to state $[m,i,n]$. If the block is added to the lower branch on some height $j\in\{d+2,\dots,m\}$, then the system transitions to state $[m,j,n]$. By Lemma 1, we have
@@
The state $[m,d,n,A]$ includes a new A-block arrival, which must be added to either the higher branch on any height up to $n+1$ or, in the case of $m<n$, the lower branch on any height up to $m+1$. Unless the block is added to the higher branch on height $n+1$ (by taking action $\overline{n+1}$ and transitioning to $[m,d,n+1]$) or the lower branch on height $m+1$ (by taking action $\underline{m+1}$ and transitioning to $[m+1,d,n]$), the system transitions to state $[m,d,n]$. In any case, all timers remain unchanged. By Lemma 1, we have
@@
5.2. Optimal Actions in the Case of $\Delta=0$
When $\Delta=0$, the private mining attack has been shown to be optimal by block-counting arguments (e.g., (guo2022bitcoin, )). Here we adopt the MDP framework to directly derive an optimal placement upon each block arrival that maximizes the probability of eventually violating the safety of height 1. While this alternative proof is not easier, the new techniques introduced here will be further developed subsequently to address the general case of $\Delta\geq 0$.
@@ 5.2. Optimal Actions in the Case of $\Delta=0$
Before proving this proposition, we first show that it is consistent with the bait-and-switch attack (Definition 15). Since $\Delta=0$, every H-blocks becomes public upon arrival. The only relevant part of Definition 15 is: “All A-blocks are kept private except when there is a unique highest branch and the height of a second highest branch is equal to the highest public height, then the second branch is made the public fork choice.” This means that if $n>m=d$, then the lower branch is made public as a “bait” to ensure that $\underline{m+1}$ is the preferred action, which is exactly rule (1) in Proposition 4. If a bait successfully attracts an H-block, subsequent H-blocks switch to this branch. Under all other conditions, bait-and-switch does not specify any particular action, except that A-blocks above the public height $d$ are nonpublic, so new H-blocks are always placed on height $d+1$, which is consistent with rule (2) in the proposition. In fact, $\underline{d+1}$ and $\overline{d+1}$ yield the same state $[m,d+1,n]$, so both are optimal under rule (2). Indeed, Definition 15 leaves this choice open. Thus, Proposition 4 establishes the optimality of bait-and-switch attack upon H-arrivals when $\Delta=0$. We relegate the proof of Proposition 4 to Appendix B.1.
@@
Here we further demonstrate that Proposition 5 for A-blocks is also consistent with the Definition 15’s description: “If there is a unique highest branch and the second-highest branch is strictly lower than the highest jumper, the adversary creates one A-block extending the second-highest branch; otherwise, the adversary creates one A-block extending a highest branch.” Evidently, this means that if $d>m$, then take action $\underline{m+1}$; otherwise, take action $\overline{n+1}$, which is consistent with Proposition 5.
@@ 6. The Case of Arbitrary Delays
In this section, we generalize the treatment of the zero-delay case in Sec. 5 to arbitrary delays upper bounded by $\Delta\geq 0$.
@@ 6.1. Preliminaries
We first justify that attack $\pi^{\prime}$ is well-defined. The only time the two systems take different actions is when an H-block is placed on height $m$ in $\mathcal{S}$, where a corresponding H-block needs to be placed on height $m+1$ in $\mathcal{S}^{\prime}$, in which case all timers in both systems remain unchanged. Hence the timer on height $m$ in $\mathcal{S}$ is always dominated by the timer on height $m+1$ in $\mathcal{S}^{\prime}$. Moreover, $\mathcal{S}$ and $\mathcal{S}^{\prime}$ always have identical timers from height $m+1$ onward. Therefore, $\pi^{\prime}$ is well-defined as long as $\pi$ is. Additionally, both branches of $\mathcal{S}^{\prime}$ are always no lower than the corresponding branches of $\mathcal{S}$ up to time $T$.
If $T$ occurs when $\mathcal{S}$’s height-$m$ timer expires, then $\mathcal{S}$ and $\mathcal{S}^{\prime}$ arrive at the same state at $T$.
@@
We next narrow down the search for an optimal action upon each new arrival.
@@
In Cases (a), since the H-block can be placed on height $j<d$, the action $\overline{d}$ is also allowable. Moreover, both $\overline{j}$ and $\underline{j}$ cause a transition to the same state as $\overline{d}$ would. Hence Case (a) is no better than $\overline{d}$.
@@
In Case (b), which is applicable only if $n>d+1$, $\overline{j}$ causes a transition to $m,j,n,(l_{j\land m},\dots,l_{d},\Delta,\dots,\Delta)]$, whereas $\overline{d+1}$ causes a transition to $m,d+1,n,(l_{(d+1)\land m},\dots,l_{d},\Delta)]$. By Lemma 8,
(36) $[m,j,n,(l_{j\land m},\dots,l_{d},\Delta,\dots,\Delta)]\prec[m,d+1,n,(l_{(d+1)\land m},\dots,l_{d},\Delta)]$
for every $j>d+1$. Hence Case (b) is no better than $\overline{d+1}$.
@@
Lemma 0 (Potentially optimal actions for an A-arrival).
  1. i)
  2. ii)
    $\underline{m+1}$ (applicable if $m<n$, causing a transition to
    $m+1,d,n,(l_{d\land(m+1)},\dots,l_{d})]$) .
Proof.
The new A-block arrival must be added to either the longer branch on any height up to $n+1$ or the lower branch on any height up to $m+1$. If $m=n$, both $\underline{m+1}$ and $\overline{n+1}$ cause a transition to the same state. Hence making $\underline{m+1}$ applicable only if $m<n$ is without loss of generality. It is easy to verify that, if either $\underline{m+1}$ (in the case of $m<n$) or $\overline{n+1}$ is taken, the corresponding new state is as described in the lemma; otherwise, the system transitions to state $S=[m,d,n,(l_{d\land m},\dots,l_{d})]$, i.e., both branches keep their heights and all timers remain unchanged. By Lemma 8, we have
(37) $S$ $\preceq[m+1,d,n,(l_{d\land(m+1)},\dots,l_{d})]$
(38) $S$ $\preceq[m,d,n+1,(l_{d\land m},\dots,l_{d})].$
This implies that either $\underline{m+1}$ or $\overline{n+1}$ is optimal. ∎
@@ 6.2. Equivalent Propositions
We establish the optimality of the bait-and-switch attack by proving the following three propositions.
@@
    In all other situations, no public fork choice is made and H-blocks are maximally delayed. Bait-and-switch would place a new H-block right above the public height, which is on height $d$ (if $l_{d}>0$) or $d+1$ (if $l_{d}=0$). In this case, Propositions 4 and 6 dictate either $\overline{d}$ or $\overline{d+1}$, depending on whether $l_{d}>0$ or not, which is consistent with bait-and-switch. In fact, $\overline{d}$ and $\underline{d}$ yield the same state, and so do $\overline{d+1}$ and $\underline{d+1}$, hence Definition 15 leaves the choice of higher or lower branch open.
@@
Due to space limitations, we relegate the rather technical proof of Propositions 46 to Appendices C.1C.2.
@@ 7. Safety at Arbitrary Target Height
We now generalize the security analysis to allow the target block to appear at an arbitrary height.
If the adversary were allowed to make unlimited attempts to mine a private fork that is $k$ blocks ahead of the public height before selecting a target block to attack, then a violation would be trivially guaranteed. To preclude such adaptive attacks, we follow Nakamoto’s original approach (nakamoto2008bitcoin, ) and impose the following constraint: An honest node generates a new public key and provides it to the sender at time $s$, who must immediately generate a transaction using the new public key and make the transaction public. This requirement ensures that the target block cannot arrive before time $s$, and will likely arrive soon after. Specifically, we assume this transaction has sufficiently high priority (e.g., a high fee rate), so that it is expected to be included in the first H-block mined after $s$. Similar non-adaptive attacks are commonly assumed to obtain safety bounds. If a target block can potentially be placed anytime in a finite execution horizon by the adversary, a union bound can be used to obtain a safety guarantee (see, e.g., (pass2017sleepy, )).
@@
Given a target transaction created at time $s$, let the block that includes this transaction be denoted as block $b$, where $h_{b}$ stands for its height, i.e., the target height. The target height is not fixed in advance—its height depends on when and where the target block is created. We are interested in the event of a safety violation occurring at the target height $h_{b}$, where block $b$ is confirmed in one of the credible chains that participates in the violation. Throughout the rest of this section and Appendix E, we use the term “safety violation of the target block” to refer to a safety violation at the target block’s height, with the target block confirmed in a credible chain.
@@
Agreement on a target block $b$ introduces an equivalence relation that partitions all blocks at height $h_{b}$ and above into two disjoint classes, depending on whether they agree on block $b$ as an ancestor. The set consisting of block $b$ and all its descendants is called the target branch. All other blocks at height $h_{b}$ and above form the challenger branch. The height of a branch is defined as the maximum height among the block(s) it contains. A branch is said to be $t$-credible if it is no lower than the public height at time $t$.
The following result is a generalization of Lemma 2 and follows from essentially the same proof given in Appendix A.1.
@@
Under the $k$-confirmation commitment rule, violating the safety of a target block $b$ is equivalent to the existence of a time $s$ at which two $s$-credible branches that disagree on block $b$ both reach height as least $h_{b}+k-1$.
@@
Under the Nakamoto consensus protocol, for a target block that is the first H-block mined after a given time $s\gg 0$, at which the pre-mining lead process has reached steady state, the bait-and-switch attack maximizes the probability of violating its safety.
@@
$1-\sum_{l=0}^{k-1}\sum_{i=0}^{k-1}\sum_{y=l}^{k-1-i}e(i)\Bigg{(}$ $f_{3}(l)\left[P^{(1)}\cdots P^{(k)}\right]_{l+1,y+1}$
(39) $+f_{4}(l)\left[P^{\prime(2)}\cdots P^{\prime(k)}\right]_{l+2,y+1}\Bigg{)}$
where $f_{3}$ is given by (8), $f_{4}$ is given by (8). Transition probability matrix $P^{(i)}$ for $i=1,\dots,k$ is given by (193)–(197). Transition probability matrix $P^{\prime(i)}$ for $i=2,\dots,k$ is given by (201)–(205).
@@
In Fig. 5, we plot the exact latency-security trade-off of Bitcoin at a general target height, given by (3) under two different adversarial fractions, with $\beta=0.25$, and $0.1$, respectively. We also include upper and lower bounds based on results in (guo2022bitcoin, ) and (doger2024refined, ).

### v2

@@ 4.6. State Comparison Rules
(25) $[m,d,n,(l_{d\land m},{\dots},l_{d}),I]\preceq[m^{\prime},d,n^{\prime},(l_{d\land m^{\prime}},\dots,l_{d}),I].$
@@
Equation (24) is true by Lemma 5. Moreover, having longer lower and higher branches improve the probability of satisfying (23), which leads to (25). ∎
@@
We can classify states based on the timer of the height just above the lower branch, $l_{m+1}$. Reaching $\mathscr{A}$ requires the lower branch to be no lower than the public height. The value of $l_{m+1}$ provides information about the race between the jumper’s height (or public height) and the lower branch height.
@@
Our categorization of states bears resemblance to the “cold”, “critical” and “hot” regions defined in (Gaži et al., 2020) in the sense that both reflect the adversary’s current advantage. The state categorization of (Gaži et al., 2020) is based on the entire blocktree and all honest views, specifically the length difference between the adversary’s best chain and the longest honest chain visible to all honest nodes. In contrast, our state categorization is based on a simplified, sufficient state representation, determined by the timer just above the second-highest branch, which allows a precise analysis of the MDP.
@@ 5. The Zero-Delay Special Case
This section is devoted to the special case of $\Delta=0$, where all honest nodes are fully synchronized at all times. An important consequence is that all H-blocks are on different heights. Also, all timers are either zero or infinity, so the public height $P=d$ in a general state representation (21). The sequence of timers in (21) always takes the form of a sequence of $\max\{1,d-m+1\}$ zeros. In this section, we omit this sequence of timers to use $[m,d,n,I]$ as a sufficient representation of the state, which is further reduced to $[m,d,n]$ when $I=\emptyset$.
@@ 5.1. Preliminaries
The state $[m,d,n,H]$ includes a new H-block arrival, which can only be added above the public height, $d$. Specifically, it must be added to either the higher branch on a height between $d+1$ and $n+1$ or the lower branch on a height between $d+1$ and $m+1$ (this is applicable only if $d\leq m$). If the block is added to the higher branch at some height $i\in\{d+2,\dots,n\}$, then the system transitions to state $[m,i,n]$. If the block is added to the lower branch at some height $j\in\{d+2,\dots,m\}$, then the system transitions to state $[m,j,n]$. By Lemma 1, we have
@@
The state $[m,d,n,A]$ includes a new A-block arrival, which must be added to either the higher branch at any height up to $n+1$ or, in the case of $m<n$, the lower branch at any height up to $m+1$. Unless the block is added to the higher branch at height $n+1$ (by taking action $\overline{n+1}$ and transitioning to $[m,d,n+1]$) or the lower branch at height $m+1$ (by taking action $\underline{m+1}$ and transitioning to $[m+1,d,n]$), the system transitions to state $[m,d,n]$. In any case, all timers remain unchanged. By Lemma 1, we have
@@
5.2. Optimal Actions in the Zero-Delay Case
When $\Delta=0$, the private mining attack has been shown to be optimal by block-counting arguments (e.g., (Guo and Ren, 2022)). Here we adopt the MDP framework to directly derive an optimal placement upon each block arrival that maximizes the probability of eventually violating the safety of height 1. While this alternative means of proof is not easier, the new techniques introduced here will be further developed subsequently to address the general case of $\Delta\geq 0$.
@@ 5.2. Optimal Actions in the Zero-Delay Case
We prove Proposition 4 in detail in Appendix B.1. The proposition is consistent with the bait-and-switch attack (Definition 15). Since $\Delta=0$, every H-blocks becomes public upon arrival. The only relevant part of Definition 15 is: “All A-blocks are kept private except when there is a unique highest branch and the height of a second highest branch is equal to the highest public height, then the second branch is made the public fork choice.” This means that if $n>m=d$, then the lower branch is made public as a “bait” to ensure that $\underline{m+1}$ is the preferred action, which is exactly rule (1) in Proposition 4. If a bait successfully attracts an H-block, subsequent H-blocks switch to this branch. Under all other conditions, bait-and-switch does not specify any particular action, except that A-blocks above the public height $d$ are nonpublic, so new H-blocks are always placed at height $d+1$, which is consistent with rule (2) in the proposition. In fact, $\underline{d+1}$ and $\overline{d+1}$ yield the same state $[m,d+1,n]$, so both are optimal under rule (2). Indeed, Definition 15 leaves this choice open. Thus, Proposition 4 establishes the optimality of bait-and-switch attack upon H-arrivals when $\Delta=0$.
@@
The proof of Proposition 5 is relegated to Appendix B.2. Here we further demonstrate that this proposition for A-blocks is also consistent with the Definition 15’s description: “If there is a unique highest branch and the second-highest branch is strictly lower than the highest jumper, the adversary creates one A-block extending the second-highest branch; otherwise, the adversary creates one A-block extending a highest branch.” Evidently, this means that if $d>m$, then take action $\underline{m+1}$; otherwise, take action $\overline{n+1}$, which is consistent with Proposition 5.
@@ 6. The Case of Arbitrary Delays
In this section, we generalize the treatment of the zero-delay case in Sec. 5 to arbitrary delays upper bounded by $\Delta\geq 0$, and henceforth develop a proof of Theorem 1.
@@ 6.1. Preliminaries
We first justify that attack $\pi^{\prime}$ is well-defined. The only time the two systems take different actions is when an H-block is placed at height $m$ in $\mathcal{S}$, where a corresponding H-block needs to be placed at height $m+1$ in $\mathcal{S}^{\prime}$, in which case all timers in both systems remain unchanged. Hence the timer at height $m$ in $\mathcal{S}$ is always dominated by the timer at height $m+1$ in $\mathcal{S}^{\prime}$. Moreover, $\mathcal{S}$ and $\mathcal{S}^{\prime}$ always have identical timers from height $m+1$ onward. Therefore, $\pi^{\prime}$ is well-defined as long as $\pi$ is. Additionally, both branches of $\mathcal{S}^{\prime}$ are always no lower than the corresponding branches of $\mathcal{S}$ up to $T$.
If the stopping time $T$ is when $\mathcal{S}$’s height-$m$ timer expires, then $\mathcal{S}$ and $\mathcal{S}^{\prime}$ arrive at the same state at $T$.
@@
We next narrow down the search for an optimal action upon each new A-arrival and H-arrival.
Lemma 0 (Potentially optimal actions for an A-arrival).
  1. i)
  2. ii)
    $\underline{m+1}$ (applicable if $m<n$, causing a transition to
    $m+1,d,n,(l_{d\land(m+1)},\dots,l_{d})]$) .
Proof.
The new A-block arrival must be added to either the longer branch at any height up to $n+1$ or the lower branch at any height up to $m+1$. If $m=n$, both $\underline{m+1}$ and $\overline{n+1}$ cause a transition to the same state. Hence making $\underline{m+1}$ applicable only if $m<n$ is without loss of generality. It is easy to verify that, if either $\underline{m+1}$ (in the case of $m<n$) or $\overline{n+1}$ is taken, the corresponding new state is as described in the lemma; otherwise, the system transitions to state $S=[m,d,n,(l_{d\land m},\dots,l_{d})]$, i.e., both branches keep their heights and all timers remain unchanged. By Lemma 8, we have
(36) $S$ $\preceq[m+1,d,n,(l_{d\land(m+1)},\dots,l_{d})]$
(37) $S$ $\preceq[m,d,n+1,(l_{d\land m},\dots,l_{d})].$
This implies that either $\underline{m+1}$ or $\overline{n+1}$ is optimal. ∎
@@
In Cases (a), since the H-block can be placed at height $j<d$, the action $\overline{d}$ is also allowable. Moreover, both $\overline{j}$ and $\underline{j}$ cause a transition to the same state as $\overline{d}$ would. Hence Case (a) is no better than $\overline{d}$.
@@
In Case (b), which is applicable only if $n>d+1$, $\overline{j}$ causes a transition to $m,j,n,(l_{j\land m},\dots,l_{d},\Delta,\dots,\Delta)]$, whereas $\overline{d+1}$ causes a transition to $m,d+1,n,(l_{(d+1)\land m},\dots,l_{d},\Delta)]$. By Lemma 8, the latter state is more favorable for every $j>d+1$. Hence Case (b) is no better than $\overline{d+1}$.
@@ 6.2. Equivalent Propositions
Building on the preliminary results in Sec. 6.1, we establish the optimality of the bait-and-switch attack by proving the following three propositions.
@@
    In all other situations, no public fork choice is made and H-blocks are maximally delayed. Bait-and-switch would place a new H-block right above the public height, which is at height $d$ (if $l_{d}>0$) or $d+1$ (if $l_{d}=0$). In this case, Propositions 4 and 6 dictate either $\overline{d}$ or $\overline{d+1}$, depending on whether $l_{d}>0$ or not, which is consistent with bait-and-switch. In fact, $\overline{d}$ and $\underline{d}$ yield the same state, and so do $\overline{d+1}$ and $\underline{d+1}$, hence Definition 15 leaves the choice of higher or lower branch open.
@@
We relegate the rather technical proof of Propositions 46 to Appendices C.1C.2.
@@ 7. Safety at Arbitrary Target Height
We now generalize the analysis to allow the target block to appear at any height. However, if the adversary could make unlimited attempts to mine a private fork that is $k$ blocks ahead of the public height before selecting a target block to attack, then a violation would be trivially guaranteed. To preclude such adaptive attacks, one may require the target block to be placed within a finite execution horizon, enabling a union bound (over all possible placements) to obtain a safety guarantee (see, e.g., (Pass and Shi, 2017b)).
In the literature, the timing of the target block is often tightly constrained. In this paper, we follow Nakamoto’s original approach (Nakamoto, 2008) and impose the following constraint on a high-value, high-priority transaction: An honest node generates a new public key and provides it to the sender at time $s$, who must immediately generate a transaction using the new public key and make the transaction public. This ensures that the target block cannot arrive before time $s$, and will likely arrive soon after. Specifically, we assume the transaction’s fee rate is high enough to ensure its inclusion in the first H-block mined after $s$.
@@
Given a target transaction created at time $s$, let the block that includes this transaction be denoted as block $b$, where $h_{b}$ stands for its height, i.e., the target height. The target height is not fixed in advance—its height depends on when and where the target block is created. We are interested in the event of a safety violation occurring at the target height $h_{b}$, where block $b$ is confirmed in one of the credible chains that participates in the violation. Throughout the rest of this section, we use the term “safety violation of the target block” to refer to a safety violation at the target block’s height, with the target block confirmed in a credible chain.
@@
Agreement on a target block $b$ introduces an equivalence relation that partitions all blocks at height $h_{b}$ and above into two disjoint classes, depending on whether they agree on chain $b$ as a prefix. The set consisting of block $b$ and all its descendants is called the target branch. All other blocks at height $h_{b}$ and above form the challenger branch. The height of a branch is defined as the maximum height among the block(s) it contains. A branch is said to be $t$-credible if it is no lower than the public height at time $t$.
The following result is a generalization of Lemma 2.
@@
Under the $k$-confirmation commitment rule, violating the safety of a target block $b$ is equivalent to the existence of a time $s$ at which two $s$-credible branches that disagree on block $b$ both reach height at least $h_{b}+k-1$.
@@
Under the Nakamoto consensus protocol, for a target block that is the first H-block mined after a given time $s$, the bait-and-switch attack maximizes the probability of violating its safety.
@@
$1-\sum_{l=0}^{k-1}\sum_{i=0}^{k-1}\sum_{y=l}^{k-1-i}e(i)$ $\Bigg{(}f_{3}(l)\left[P^{(1)}\times\cdots\times P^{(k)}\right]_{l+1,y+1}$
(38) $+f_{4}(l)\left[P^{\prime(2)}\times\cdots\times P^{\prime(k)}\right]_{l+2,y+1}\Bigg{)}$
where $f_{3}$ is given by (8), $f_{4}$ is given by (8), the transition probability matrix $P^{(i)}$ for $i=1,\dots,k$ is given by (192)–(196), and the transition probability matrix $P^{\prime(i)}$ for $i=2,\dots,k$ is given by (200)–(204).
@@
In Fig. 7, we plot the exact latency-security trade-off of Bitcoin at a general target height, given by (3) under two different adversarial fractions, with $\beta=0.25$, and $0.1$, respectively. We also include upper and lower bounds based on results in (Guo and Ren, 2022) and (Doger and Ulukus, 2025a).